# Masked Training for Robust Arrhythmia Detection from Digitalized Multiple Layout ECG Images

### v1

@@ METHODS
In our designed model, each block has two inherent features, different leads and different time positions. Unlike simple time series, we need to know the relative position of each block in the ECG. In order to enable PatchECG to better capture the relative position of each patch in the ECG, we initialized a lead embedding list $LE=\left\{l_{0},l_{1},\dots,l_{C}\right\}$ and a time embedding list $TE=\left\{t_{0},t_{1},\dots,t_{\left\lfloor\frac{C}{P}\right\rfloor}\right\}$, Both can be learned during the training process. Add the lead code and time code to $\bar{m}$ to obtain:
$\bar{m}=\left\{M_{i,j}+l_{i}+t_{j}\mid i=0,1,\dots C,j=0,1,\dots\left\lfloor\frac{C}{P}\right\rfloor\right\}$ (4)
Discard completely missing patches and expand the remaining patches to obtain $\bar{X}\in\mathbb{R}^{(N-N_{miss})\times D}$, where $N_{miss}$ represents the number of completely missing patches. By optimizing and rearranging $\bar{X}$ to form a new $\bar{X^{\prime}}$ , potential spatiotemporal dependencies can be better captured. Firstly, segment $\bar{X}$ into multiple segments, then shuffle these segments to optimize the original and shuffled patch sequences through a weighted sum of learnable weights in the network. Finally, use the stitch module to connect the shuffled connections and output them to the next layer. Stack them in the neural network to obtain the output $\bar{X^{\prime}}$ 24.
@@
All implementation details match the baseline. We use the exact hyperparameters of the baseline specified in the original paper, or when the code is available. Alternatively, when hyperparameters are not precisely specified in the paper or code, we attempt to maximize performance by searching for the best hyperparameters ourselves. During the training process of PatchECG, the window size $P$ is set to 64, the batch size is set to 64, and the number of training epochs is 30. The optimizer uses Adam, with an initial learning rate of $1\times 10^{-3}$ and a weight decay coefficient of $1\times 10^{-4}$. In the PatchEncoder module, the number of basic filters is set to 16, the size of the convolution kernel is 16, and the number of filters in each layer is [16, 32, 32, 40, 40, 64, 64]. Each layer contains two convolution blocks. The hidden feature dimension of the Transformer module is set to 768, with a depth of 3 layers and an 8-head attention mechanism. The S324 module consists of 3 layers, with an initial number of 4 segments and a segment multiplier of 2. Our code is implemented using PyTorch, and our experiments were conducted on NVIDIA RTX 3090Ti GPU.
@@
  • The dataset used 21388 clinical 12-lead ECGs from 18869 patients using PTB-XL, with each sample lasting 10 seconds. 52% are males and 48% are females, with an age range of 0-95 years old 25. We use multi label classification for this data: diagnostic subclasses. Specifically, the experiment used the official 10 fold split scheme provided by the PTB-XL dataset: in the first round, folds 1-8 were used as the training set, fold 9 as the validation set, and fold 10 as the test set; In the second round, the 7th fold is used as the validation set, the 8th fold is used as the test set, and the remaining folds are used as the training set; Repeat this process for the remaining rounds, rotating the division of the validation and testing sets to complete the evaluation of all folds. Use the Subclass (23 classes) of the multi-class classification task as the label.
@@
  • We used the tool provided by ECG-image-kit to generate 21388 ECG images with 3 x 4 layouts of different noise levels on the PTB-XL dataset, and further transformed these images into high-quality ECG signal data that can be used for subsequent analysis using the digital tool 5, which ranked first in the 2024 CINC competition.
@@
  • To evaluate the generalization performance of the model, this study selected a cohort of 400 AF patients admitted to Chaoyang Hospital as an independent external validation set for performance verification. Due to differences in image resolution, size, and layout among the ECG data provided by Chaoyang Hospital, it is difficult to effectively process them using existing automated tools. Therefore, we used Paper ECG 6 to manually digitize the image data to ensure the accuracy of the signal data.
@@
In our experiment, the task type is the multi-class classification. Specifically, during the model training and evaluation phase, we use the average Area Under the Receiver Operating Characteristic Curve (Avg AUROC) as the main evaluation metric; When conducting external validation on the AF dataset at Chaoyang Hospital, we comprehensively considered multiple indicators such as Accuracy, Precision, Recall, Specification, F1, and AUROC, with a particular focus on recall and specificity performance. In addition, due to significant label imbalance issues in the dataset used, we introduced the Focal Loss loss function 26 during the training process: $loss\left(p\right)=-\alpha_{t}{\left(1-p\right)}^{\gamma}\log{p}$ To enhance attention to categories with fewer quantities and alleviate the adverse effects of category imbalance.
RESOURCE AVAILABILITY
ACKNOWLEDGMENTS
This work was supported by the National Natural Science Foundation of China (62102008); CCF-Zhipu Large Model Innovation Fund (CCF-Zhipu202414).
DECLARATION OF INTERESTS

### v3

@@ Methods
In our designed model, each block has two inherent features, different leads and different time positions. Unlike simple time series, we need to know the relative position of each block in the ECG. In order to enable PatchECG to better capture the relative position of each patch in the ECG, we initialized a lead embedding list $LE=\left\{l_{0},l_{1},\dots,l_{C}\right\}$ and a time embedding list $TE=\left\{t_{0},t_{1},\dots,t_{\left\lfloor\frac{c_{i}}{P}\right\rfloor}\right\}$, Both can be learned during the training process. Add the lead code and time code to $\bar{m}$ to obtain:
$\bar{m}=\left\{M_{i,j}+l_{i}+t_{j}\mid i=0,1,\dots C,j=0,1,\dots\left\lfloor\frac{c_{i}}{P}\right\rfloor\right\}$ (5)
To better capture the collaborative dependencies between non-adjacent patches across leads and time, we adopt the Segment, Shuffle, and Stitch (S3) module 11 to rearrange the patch sequence $\bar{X}\in\mathbb{R}^{(N-N_{\text{miss}})\times D}$ in a task-driven manner.
Segment. The input sequence $\bar{X}$ is divided into $n_{s}$ non-overlapping segments of equal length $\tau_{s}=\lfloor(N-N_{\text{miss}})/n_{s}\rfloor$, yielding a segment set $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n_{s}}\}\in\mathbb{R}^{\tau_{s}\times D\times n_{s}}$, where $s_{k}=\bar{X}[(k-1)\tau_{s}:k\tau_{s}]$. The initial number of segments is set to $n_{s}=4$, following the ablation study in the original S3 work 11, which demonstrates that beginning with a coarse-grained partition and progressively refining it achieves a better trade-off between capturing global inter-lead dependencies and local temporal coherence. When $(N-N_{\text{miss}})$ is not exactly divisible by $n_{s}$, the first $(N-N_{\text{miss}})\bmod n_{s}$ patches are truncated before segmentation and appended back to the output of the final S3 layer, ensuring no patch information is discarded.
Shuffle. A learnable shuffling vector $\mathbf{Q}=\{q_{1},q_{2},\dots,q_{n_{s}}\}\in\mathbb{R}^{n_{s}}$ determines the optimal reordering of segments. Each $q_{k}$ corresponds to segment $s_{k}$; segments are reordered according to the descending rank of $\mathbf{Q}$:
$\mathcal{S}^{\text{shuf}}=\text{Sort}(\mathcal{S},\ \text{key}=\mathbf{Q})$ (6)
To maintain differentiability, the permutation is implemented via a binary assignment matrix $\tilde{\Omega}$ derived from $\sigma_{\text{idx}}=\text{Argsort}(\mathbf{Q})$, allowing gradients to flow through $\mathbf{Q}$ during backpropagation 11. All elements of $\mathbf{Q}$ are initialized to the same constant value of $0.6$, so that the initial shuffling corresponds to the identity permutation; the optimal reordering is then learned jointly with the rest of the network in a task-driven manner.
Stitch. The shuffled segments are concatenated to reconstruct a full-length sequence $X^{\text{shuf}}\in\mathbb{R}^{(N-N_{\text{miss}})\times D}$. To preserve the information in the original ordering while incorporating the newly learned arrangement, a learned weighted sum is performed via a Conv1D layer with learnable weights $w_{1}$ and $w_{2}$:
$\bar{X}^{\text{out}}=w_{1}\cdot\bar{X}+w_{2}\cdot X^{\text{shuf}}$ (7)
Both $w_{1}$ and $w_{2}$ are initialized to $1$ and updated jointly with all model parameters throughout training.
Stacking. The S3 module is stacked for $\varphi=3$ layers to enable multi-granularity shuffling. In each subsequent layer $\ell$, the number of segments grows according to:
$n_{s}^{(\ell)}=n_{s}\times\theta^{\ell-1},\quad\ell=1,2,\dots,\varphi$ (8)
where $\theta=2$ is the segment multiplier, yielding $n_{s}^{(1)}=4$, $n_{s}^{(2)}=8$, and $n_{s}^{(3)}=16$ segments across the three layers. This progressive refinement allows the model to first capture coarse inter-lead dependencies at the block level, then refine to finer-grained intra-lead temporal structures. The output of the final S3 layer is $\bar{X}^{\prime}\in\mathbb{R}^{(N-N_{\text{miss}})\times D}$, which encodes both the original and task-optimally reordered patch representations.
@@
All implementation details match the baseline. We use the exact hyperparameters of the baseline specified in the original paper, or when the code is available. Alternatively, when hyperparameters are not precisely specified in the paper or code, we attempt to maximize performance by searching for the best hyperparameters ourselves. During the training process of PatchECG, the window size $P$ is set to 64, the batch size is set to 64, and the number of training epochs is 30. The optimizer uses Adam, with an initial learning rate of $1\times 10^{-3}$ and a weight decay coefficient of $1\times 10^{-4}$. In the PatchEncoder module, the number of basic filters is set to 16, the size of the convolution kernel is 16, and the number of filters in each layer is [16, 32, 32, 40, 40, 64, 64]. Each layer contains two convolution blocks. The hidden feature dimension of the Transformer module is set to 768, with a depth of 3 layers and an 8-head attention mechanism. The S311 module consists of 3 layers, with an initial number of 4 segments and a segment multiplier of 2. Our code is implemented using PyTorch, and our experiments were conducted on NVIDIA RTX 3090Ti GPU.
Table 1: Model Parameters and Inference Time Comparison.
Model Parameters (M) Inference Time (ms)
PatchECG 21.72 1.094
ECG-Founder 30.82 23.900
Medformer 12.11 1.350
SimMTM 12.20 0.021
SimMTM-KNN 12.20 4.872
SimMTM-SAITS 13.55 1.319
TimeXer 13.63 0.222
  • Note: Bold values indicate the best performance in each column.
To assess the practical deployment feasibility of PatchECG, we report the model parameter count and per-sample inference time for all compared methods in Table 1. As shown, PatchECG contains 21.72M parameters and achieves an inference time of 1.094 ms per sample, which is comparable to Medformer and SimMTM-SAITS, and substantially faster than ECGFounder. Although SimMTM has the shortest inference time (0.021 ms), it achieves substantially lower AUROC across all layouts, indicating that its speed advantage comes at a significant cost to diagnostic accuracy. These results demonstrate that PatchECG strikes a favorable balance between model complexity, inference efficiency, and diagnostic performance, making it suitable for real-time clinical deployment scenarios.
@@
  • The dataset used 21388 clinical 12-lead ECGs from 18869 patients using PTB-XL, with each sample lasting 10 seconds. 52% are males and 48% are females, with an age range of 0-95 years old 28. We use multi label classification for this data: diagnostic subclasses. Specifically, the experiment used the official 10 fold split scheme provided by the PTB-XL dataset: in the first round, folds 1-8 were used as the training set, fold 9 as the validation set, and fold 10 as the test set; In the second round, the 7th fold is used as the validation set, the 8th fold is used as the test set, and the remaining folds are used as the training set; Repeat this process for the remaining rounds, rotating the division of the validation and testing sets to complete the evaluation of all folds. Use the Subclass (23 classes) of the multi-class classification task as the label.
@@
  • We used the tool provided by ECG-image-kit to generate 21388 ECG images with 3 x 4 layouts of different noise levels on the PTB-XL dataset, and further transformed these images into high-quality ECG signal data that can be used for subsequent analysis using the digital tool 17, which ranked first in the 2024 CINC competition.
@@
  • To evaluate the generalization performance of the model, this study selected a cohort of 400 AF patients admitted to Chaoyang Hospital as an independent external validation set for performance verification27, 26. The clinical characteristics of this cohort are summarized in Table 2. Due to differences in image resolution, size, and layout among the ECG data provided by Chaoyang Hospital, it is difficult to effectively process them using existing automated tools. Therefore, we used Paper ECG 8 to manually digitize the image data to ensure the accuracy of the signal data.
Table 2: Clinical characteristics of the Chaoyang Hospital cohort.
Characteristics Chaoyang cohort ($N=400$)
Age (years) 67.1 $\pm$ 10.3
Sex
   Male 189 (47.25%)
   Female 211 (52.75%)
BMI (kg/m2) 26.2 $\pm$ 3.7
Hypertension 265 (66.25%)
Diabetes 92 (23.00%)
Heart failure 17 (4.25%)
Smoking status 1 (0.25%)
Stroke 1 (0.25%)
Atrial Fibrillation 200 (50%)
Follow-up time (years) 4.82 $\pm$ 3.79
  • Continuous variables (e.g., age and BMI) are presented as mean $\pm$ standard deviation. Categorical variables are presented as count (percentage).
Since the Chaoyang Hospital dataset contains only ECG image files without corresponding ground-truth digital signals, a direct evaluation of the Paper ECG digitization quality on this cohort is not feasible. To provide an indirect assessment of digitization accuracy, we conducted a controlled comparison using a separate subset of ECG images for which reference signals were available.
[…]
$\text{SNR}=10\log_{10}\frac{\sum_{t}x(t)^{2}}{\sum_{t}[x(t)-\hat{x}(t)]^{2}}$ (14)
$\text{PSNR}=10\log_{10}\frac{x_{\max}^{2}}{\frac{1}{N}\sum_{t}[x(t)-\hat{x}(t)]^{2}}$ (15)
where $x(t)$ denotes the reference signal, $\hat{x}(t)$ the digitized signal, $x_{\max}$ the peak amplitude of the reference, and $N$ the total number of samples. Higher values of both SNR and PSNR indicate better digitization fidelity.
The results yielded an Avg SNR of $\bm{-1.59}$ dB and an Avg PSNR of $\bm{12.52}$ dB. We note that the relatively low numerical values do not necessarily reflect poor digitization quality, for two reasons. First, the reference signals used in this comparison are themselves derived from an independent digitization pipeline rather than ground-truth electronic recordings, introducing an inherent ceiling on achievable agreement. Second, minor temporal misalignments between the two digitized signals — even at the sub-sample level — can substantially penalize both SNR and PSNR due to their point-wise sensitivity to phase offsets. Visual inspection of the digitized waveforms confirms that the Paper ECG outputs closely follow the reference signals, with nearly identical morphology across all major waveform components, as illustrated in Figure 4.
Figure 4: Overlay of original and digitized ECG signals.
Figure 4 shows representative examples of the digitization quality comparison: panel Left presents a single I Lead segment of 2.5 seconds, and panel Right presents a complete II lead recording. In both panels, the blue trace represents the signal digitized by Paper ECG and the red trace represents the output of our in-house algorithm. The two signals exhibit strong visual agreement, supporting the conclusion that the Paper ECG tool produces reliable digitization output suitable for downstream model evaluation.
@@
In our experiment, the task type is the multi-class classification. Specifically, during the model training and evaluation phase, we use the average Area Under the Receiver Operating Characteristic Curve (Avg AUROC) as the main evaluation metric; When conducting external validation on the AF dataset at Chaoyang Hospital, we comprehensively considered multiple indicators such as Accuracy, Precision, Recall, Specification, F1, and AUROC, with a particular focus on recall and specificity performance. In addition, due to significant label imbalance issues in the dataset used, we introduced the Focal Loss loss function 19 during the training process: $loss\left(p\right)=-\alpha_{t}{\left(1-p\right)}^{\gamma}\log{p}$ To enhance attention to categories with fewer quantities and alleviate the adverse effects of category imbalance.
Results
To comprehensively evaluate the performance of our proposed PatchECG, we conducted a series of rigorous experiments and compared it with several key baseline models, including: 1) The typical baseline 22 in the main follow MissTSM paper uses zero padding, KNN interpolation, and SAITS interpolation6, respectively, and combines them with SimMTM for prediction 5; 2) Medformer is a model specifically designed for biosignals, which introduces six different noise perturbations intra- and inter-granularity during the training phase, thus enhancing robustness to missing 31;3) TimeXer, on the other hand, provides auxiliary cues to the model by explicitly modeling missing patterns and treating the missing as a kind of external information32. We explored the diagnostic accuracy of the generated signal data digitized from ECG images by different layouts of signal data and the signal data digitized from different layouts of ECG images from the real Chaoyang Hospital.
PatchECG Achieve Better Results on Different Layouts
Figure 5: Results of PatchECG and Other Methods with Different Layouts. Box plots represent the average AUROC performance across 5-fold cross-validation ($n=2158$ ECG samples per test fold). For each box plot, the center line indicates the median, the box limits indicate the upper and lower quartiles (interquartile range), and the whiskers extend to the minimum and maximum values.
PatchECG is Highly Scalable
To effectively utilize existing excellent models, we explored three different modules for encoding patches. We primarily adopted representative models such as Net1D, ECGFounder, and Projection as PatchEncoder:
  • Projection: Inspired by PatchTST 23, it maps each patch to the embedding space through a single fully connected layer. This module has the simplest structure with no temporal feature extraction capability.
  • Net1D: A residual one-dimensional convolutional model combined with an SE attention module 12, trained from scratch on the target dataset. It is capable of learning local temporal features and channel-wise dependencies within each patch.
  • ECGFounder: The latest large-scale ECG foundation model pre-trained on 10 million ECGs 18. Crucially, ECGFounder shares an identical backbone architecture with Net1D; the sole difference is that its parameters are frozen and directly transferred from large-scale pre-training, rather than trained from scratch on the target dataset.
Table 3: Different modules as PatchEncoder
Method
3 x 4
3 x 4 + II
3 x 4 + II + V1
6 x 2
6 x 2 + II
12 x 1
Random
Ours-Net1D 0.831 $\pm$ 0.019 0.829 $\pm$ 0.019 0.830 $\pm$ 0.022 0.833 $\pm$ 0.022 0.832 $\pm$ 0.022 0.829 $\pm$ 0.026 0.835 $\pm$ 0.022
  • Note: Bold values indicate the best performance in each column.
We adopted the same training, validation, and test set partitions as in Experiment 1 and conducted 5-fold cross-validation. As shown in Table 3, all three PatchEncoder variants maintain consistent AUROC across different layouts, confirming that PatchECG effectively captures asynchronous signal dependencies regardless of layout configuration. A detailed analysis of each module is as follows. Projection performs the worst due to its simple linear structure, which lacks the capacity to extract temporal features from patches. Net1D is trained from scratch and achieves the second-best result, demonstrating that a residual convolutional structure with SE attention can effectively learn patch-level representations. ECGFounder achieves the best average AUROC across all layouts. Notably, ECGFounder shares the same backbone architecture as Net1D; the sole difference is that its parameters are frozen weights transferred from pre-training on 10 million ECGs, rather than trained from scratch. This architectural equivalence enables a direct comparison: the consistent performance gain of approximately 0.04–0.05 AUROC over Net1D can be attributed specifically to the rich ECG morphology priors encoded by large-scale pre-training, rather than to any structural advantage. In addition, freezing the pre-trained parameters helps prevent overfitting, resulting in more stable performance across layouts.
In summary, PatchECG can effectively incorporate existing models as PatchEncoders, including large-scale pre-trained foundations. Combined with the adaptive variable block number missing mechanism, PatchECG maintains high accuracy across both asynchronous and synchronous ECG data with different layouts.
PatchECG Achieve Better Results on Generating ECG Images
[…]
Figure 6: Digitally generated ECG image. Box plots represent the average AUROC performance evaluated through 5-fold cross-validation ($n=21388$ digitally generated ECG samples). For each box plot, the center line indicates the median, the box limits indicate the upper and lower quartiles (interquartile range), and the whiskers extend to the minimum and maximum values.
[…]
PatchECG Achieve Better Results on Real Hospital ECG Images
On the external test set of Chaoyang Hospital, which mainly consists of 12 x 1 layout, with some being 3 x 4 and 6 x 2 layout (as shown in Figure 7), our method achieved the best AUROC among all AF baseline predictions (as shown in Table 4). The overall performance of ECGFounder’s predictions is good. Although some methods have higher recall rates, their specificity is significantly lower, making them prone to misclassifying negative samples as positive, leading to an increased risk of misdiagnosis. Due to pre-training on tens of millions of ECGs, ECGFounder exhibits strong robustness. When there is a certain degree of data loss, the method fills in the data to a complete 12 x 1 layout through zero-padding before inputting it into the model. On the external AF cohort of Chaoyang Hospital, the three imputation strategies performed similarly in terms of AUROC scores, with almost no significant differences. This indicates that imputation does not significantly improve the model’s accuracy. Medformer performed the best among all baseline methods. This method introduces six data augmentation strategies during the training stage, which act on its internal and external granularity modeling modules, respectively, to enhance the model’s robustness to noise and missing data. When dealing with the Chaoyang Hospital cohort with different layouts and poor digitization quality, Medformer achieved an AUROC of 0.605, demonstrating strong adaptability and stability. TimeXer performed well on high-quality digitized signal data, but its performance significantly decreased in the Chaoyang Hospital cohort with poor digitization quality. This method treats missing data as exogenous information and uses it to assist in the modeling and prediction of endogenous information. However, when digitization is poor, not only is the endogenous information itself inaccurate, but the generated missing pattern information also becomes inaccurate. The exogenous information originally used to guide the model becomes a source of interference instead. The introduction of this noise weakens the model’s judgment ability, leading to unsatisfactory prediction results on real Chaoyang Hospital image data.
Method Accuracy Precision Recall Specificity F1 AUROC
Ours 0.745 0.743 0.750 0.740 0.746 0.778
ECGFounder $0.625$ $0.605$ $0.720$ $0.530$ $0.658$ $0.667$
SimMTM $0.463$ $0.332$ $0.600$ $0.395$ $0.427$ $0.541$
SimMTM-KNN $0.485$ $0.480$ $0.540$ $0.430$ $0.512$ $0.539$
SimMTM-SAITS $0.490$ $0.491$ $0.550$ $0.430$ $0.519$ $0.541$
Medformer $0.498$ $0.498$ $0.640$ $0.335$ $0.560$ $0.605$
TimeXer $0.489$ $0.487$ $0.460$ $0.515$ $0.473$ $0.485$
  • Note: Bold values indicate the best performance in each column.
Method Accuracy Precision Recall Specificity F1 AUROC
Ours 0.833 0.723 $0.810$ 0.845 0.764 0.893
ECGFounder $0.573$ $0.430$ 0.860 $0.430$ $0.573$ $0.703$
SimMTM $0.485$ $0.486$ $0.525$ $0.445$ $0.505$ $0.450$
SimMTM-KNN
SimMTM-SAITS
Medformer $0.450$ $0.332$ $0.640$ $0.355$ $0.437$ $0.510$
TimeXer $0.473$ $0.289$ $0.390$ $0.510$ $0.331$ $0.423$
  • Note: Bold values indicate the best performance in each column.
[…]
It should be noted that none of the models in this study, including PatchECG, underwent threshold recalibration on the Chaoyang Hospital data, ensuring a fair comparison across all methods. Since recall and specificity are both threshold-dependent, the AUROC metric provides a more appropriate measure of intrinsic discriminative ability. ECGFounder’s AUROC of 0.703 is substantially lower than PatchECG’s 0.893, confirming that the gap reflects a genuine difference in model capability rather than a threshold artifact.
Case Study
Quantitative Evaluation of Model Interpretability
To quantitatively assess whether the model’s attention focuses on clinically relevant ECG segments, we conducted a comparison between the model’s patch-level attention scores and annotations from two experienced cardiologists. Specifically, 20 ECG samples (10 AF and 10 NAF) were selected from the Chaoyang Hospital cohort. Each cardiologist independently identified the top 20 patches they considered most diagnostically relevant. The model’s top 20 patches were determined by ranking the patch attention scores in descending order. Overlap was evaluated using two complementary metrics:
Top-$K$ Overlap Rate (Top-$K$ OR), defined as the proportion of patches shared between the model’s top-$K$ set and the clinician’s top-$K$ set:
$\text{Top-}K\text{ OR}=\frac{|\mathcal{M}_{K}\cap\mathcal{C}_{K}|}{K}\times 100\%$ (16)
Jaccard Similarity Index (JSI), defined as:
$\text{JSI}=\frac{|\mathcal{M}_{K}\cap\mathcal{C}_{K}|}{|\mathcal{M}_{K}\cup\mathcal{C}_{K}|}\times 100\%$ (17)
where $\mathcal{M}_{K}$ and $\mathcal{C}_{K}$ denote the top-$K$ patch sets of the model and the clinician, respectively, and $K=20$ in all experiments. Higher values indicate greater agreement.
The results are summarized in Table 6. The inter-rater agreement between the two cardiologists serves as an upper-bound reference, yielding a Top-$K$ OR of 41.8% $\pm$ 15.8% and a JSI of 27.8% $\pm$ 13.9%, reflecting the inherent variability in clinical judgment even among experienced physicians. The model’s agreement with Doctor B reaches a Top-$K$ OR of 36.0% $\pm$ 11.9% and a JSI of 22.6% $\pm$ 9.2%, approaching the inter-rater ceiling. The lower agreement with Doctor A (Top-$K$ OR: 20.5% $\pm$ 10.7%; JSI: 11.8% $\pm$ 6.8%) is consistent with the naturally high inter-clinician variability observed in patch-level ECG annotation tasks. These results collectively suggest that the model’s attention mechanism captures diagnostically meaningful ECG segments that are broadly consistent with clinical expert judgment.
Table 6: Quantitative comparison between model and cardiologist on ECG ($K=20$).
Comparison Top-$K$ OR (%) JSI (%)
Model vs Doctor A $20.5\pm 10.7$ $11.8\pm 6.8$
Model vs Doctor B $36.0\pm 11.9$ $22.6\pm 9.2$
Doctor A vs Doctor B $41.8\pm 15.8$ $27.8\pm 13.9$
Discussion
[…]
Regarding data usage and experimental setup, we use PTB-XL data as the training data source in this study. There are certain limitations in data usage, which to some extent restrict the model’s generalization ability to broader patient populations and more complex pathological types. During the testing phase, we utilize the ECG-image-kit tool to generate ECG images and digitize them to evaluate the model. However, due to limitations in digitization methods, only 3 x 4 layout can be accurately processed, and other layout types of images cannot be effectively digitized. Therefore, in the generated images, we primarily focus on the 3 x 4 layout. In recent years, various digitization methods have emerged, but solutions that truly efficiently and accurately restore high-quality signals remain limited. It is particularly noteworthy that current digitization methods generally lack effective mechanisms to deal with lead adhesion. When lead adhesion occurs, the signals in the image overlap, which can easily be misjudged by the model as pathological features such as ST segment deviation, leading to deviations in prediction results. Therefore, how to improve the digitization quality of ECG images, especially developing digitization algorithms capable of identifying and correcting lead adhesion issues, is one of the important directions for achieving higher-quality AI-ECG diagnosis research in the future.
Conclusion
Data availability
The original PTB-XL dataset used in this study is publicly available on PhysioNet at https://physionet.org/content/ptb-xl/1.0.3/. The derivative dataset generated and analyzed during the current study—comprising 1-dimensional ECG signals reverse-digitized from synthesized multiple-layout ECG images (using the ecg-image-kit and the 2024 Computing in Cardiology Challenge 1st place digitization algorithm)—is publicly available in the Zenodo repository at https://zenodo.org/records/19254077. The diagnostic labels for these digitized signals correspond directly to the ground-truth annotations of the original PTB-XL dataset.The Beijing Chaoyang Hospital Institutional Research Board approved the study protocol and all ethical considerations (No. 2021-Ke-555). The need for informed consent was waived by the ethics board of our hospital because the images had been acquired during daily clinical practice.
Code availability
The imolementation of PatchECG is available in the GitHub repositorv at https://github.com/zhangshanwei-mm/PatchECG. PaperECG is available in the GitHub repository at https://github.com/Tereshchenkolab/paper-ecg/tree/master. Our internal digitization algorithms are available in the GitHub repository at https://github.com/PKUDigitalHealth/ecg-img2ts.The open-source ecg-image-kit tool utilized for generating ECG images is available a https://github.com/alphanumericslab/ecg-image-kit. The automated digitization of ECG image into digital signals was performed using ECG Digitiser, the 1st-place automated digitization algorithm in the George B. Moody PhysioNet Challenge 2024. The code is available at https://github.com/felixkrones/ECG-Digitiser.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant 62102008 to SH; Grant 62376197, 62020106004 and 92048301 to YZ); CCF - Tencent Rhino-Bird Open Research Fund (Grant CCF-Tencent RAGR20250108 to SH); CCF-Zhipu Large Model Innovation Fund (Grant CCF-Zhipu202414 to SH); PKU-OPPO Fund (Grant BO202301 and BO202503 to SH);Research Project of Peking University in the State Key Laboratory of Vascular Homeostasis and Remodeling (Grant 2025-SKLVHR-YCTS-02 to SH); Beijing Municipal Science and Technology Commission (Grant Z251100000725008 to SH); Tianchi Elite Youth Doctoral Program (Grant CZ002701 and CZ002707 to YZ).
Author information
Authors and Affiliations
Department of Computer Science, Tianjin University of Technology, Tianjin, China
Shanwei Zhang, Kexin Wang, Yuxi Zhou & Shengyong Chen
HeartVoice Medical Technology, Hefei, China
Deyun Zhang & Shijia Geng
Department of Cardiology, Beijing Chaoyang Hospital, Capital Medical University, Beijing, China
Yirao Tao & Xingpeng Liu
Department of Cardiology, Peking University People’s Hospital, Beijing, China
Qinghao Zhao
DCST, BNRist, RIIT, Institute of Internet Industry, Tsinghua University, Beijing, China
Yuxi Zhou
Tianjin Key Laboratory of Ionic-Molecular Function of Cardiovascular Disease, Department of Cardiology, Tianjin Institute of Cardiology, the Second Hospital of Tianjin Medical University, Tianjin 300211, China
Xingliang Wu
National Institute of Health Data Science, Peking University, Beijing, China
Shanwei Zhang, Jun Li & Shenda Hong
Institute for Artificial Intelligence, Peking University, Beijing, Chin
Shenda Hong
Institute of Medical Technology, Peking University Health Science Center, Beijing, China
Shenda Hong
State Key Laboratory of Vascular Homeostasis and Remodeling, NHC Key Laboratory of Cardiovascular Molecular Biology and Regulatory Peptides, Peking University, Beijing, China
Shenda Hong
Corresponding author
Correspondence to Shenda Hong & Yuxi Zhou.
Contributions
S.Z. performed the primary data processing, algorithm implementation, and drafted the initial version of the manuscript.Q.Z. and X.W. provided clinical expertise and conducted the medical evaluation of the model’s interpretability. D.Z., Y.T., and K.W. contributed to data aggregation, baseline preparation. S.G., J.L., and X.L. assisted with the design of the validation framework and carried out additional data analyses. S.C. contributed to the conceptualization of the methodology and co-developed the proposed solution through critical discussions. Y.Z. and S.H. conceived and supervised the overall study, provided critical revisions to the manuscript, and secured access to data and computational resources. All authors critically reviewed and approved the final manuscript.
Ethics declarations
Competing interests